OpenStudy (anonymous):

x-4=4 how can I make 3 ordered pairs ?

OpenStudy (saifoo.khan):

3 pairs?

OpenStudy (saifoo.khan):

Impossible.

OpenStudy (anonymous):

lol your good dude well i screwed up so x-4=4

OpenStudy (saifoo.khan):

lol. so x = 4 + 4 x = 8

OpenStudy (radar):

That does have a single solution.

jimthompson5910 (jim_thompson5910):

You're probably missing a y term in there somewhere.

OpenStudy (saifoo.khan):

as radar said, it srsly have a single solution

OpenStudy (anonymous):

it said 3 pairs

OpenStudy (saifoo.khan):

might be a mistake.

OpenStudy (radar):

Does 3 pairs beat a straight? lol

OpenStudy (saifoo.khan):

haha @ radar.

OpenStudy (saifoo.khan):

hi myinin

myininaya (myininaya):

we have x=0 so you having the following pairs on that line: (0,-5) (0,0) (0,1/2) (0,3) every point on the y axis is on that line

myininaya (myininaya):

oops x=8

myininaya (myininaya):

some ordered pairs are (8,-5) (8,1/2) (8,0) (8,3) and so on...

myininaya (myininaya):

on in general (8,y) is on that line

myininaya (myininaya):

hey saifoo

OpenStudy (anonymous):

zapos

myininaya (myininaya):

satellite i threw a party for you and you never noticed ask saifoo

OpenStudy (anonymous):

actually it is zappos

OpenStudy (anonymous):

when? where?

OpenStudy (saifoo.khan):

Yeah..

myininaya (myininaya):

its too late i will look it up and give you the link